You have a total of $5,000 and decide to invest part in a bank paying 5% simple interest and in a CD that pays 8%. If you receive a total of $300 from your investments, how much was put in each account?
Answer by checkley75(3666)

.08x+.05(5,000-x)=300
.08x+250-.05x=300
.03x=300-250
.03x=50
x=50/.03
x=1,666.67 invested @ 8%
5,000-1,666.67=3,333.33 invested @ 5%
Proof:
.08*1,666.67+.05*3,333.33=300
133.33+166.67=300
300=300